find 3 consecutive odd numbers where the product of the smaller two numbers is 46 less than the square of the largest number
Answer by stanbon(75887)

find 3 consecutive odd numbers where the product of the smaller two numbers is 46 less than the square of the largest number
--------------
1st: x-1
2nd: x
3rd: x+1
-------------------
Equation:
x(x-1) = (x+1)^2-46
---
x^2-x = x^2+2x-45
-x = 2x-45
-3x = -45
x = 15
x-1 = 14
x+1 = 16
--------------
Cheers,
stan H.